# Convergent transformations of visual representation in brains and models

### v1

@@ A.6 Power-Law Attenuation Model of RSA Metrics
Figure A8: Global power-law model captures every pair-wise relation among RSA metrics. Scatter plots (grey points) show parcel-wise observed RSA values for all metric combinations: inter-subject (shifted and unshifted), within-subject, subject-versus-group, and vision-model correlations, each additionally split by hemisphere sampling (left, right, both). Red curves depict an independent power-law fit for the specific scatter, whereas black dashed curves are the predictions from a single joint fit of Eq. (A3) applied to all metrics simultaneously (no panel–specific re–fitting). The accompanying table reports the fitted scale factors $a_{j}$, exponents $b_{j}$, and explained variance $R^{2}$ for every metric (model-predicted vs observed rsa). All metric trends are explained with $R^{2}\!>\!0.92$, confirming that one latent similarity map, together with metric-specific attenuation parameters, suffices to account for the entire measurement set.
B THINGS-fMRI and BOLD5000 Processing
B.1 THINGS-fMRI dataset
@@
We processed the THINGS-fMRI dataset 25 using a pipeline analogous to that for NSD. This dataset comprises 3T fMRI recordings from three participants performing an oddball detection task on 8,740 unique object images. A subset of 100 images was repeated once per session (12 presentations total); all other images appeared only once. For our primary computations, we used only the first presentation of each of the 8,740 images.
@@
$\mathcal{A}_{p,q,r,r^{\prime},s}^{\mathrm{IS}}=\mathrm{RSA}\bigl{(}X_{p,r,s},X_{q,r^{\prime},s}\bigr{)}.$ (B4)
@@
Group-level maps were then obtained by averaging these scores across sessions and all participant pairs. Model-brain RSA followed the same session-wise procedure. As a control, we confirmed that computing IS-RSA over the full 8,740-trial set (using $,740\times 8$ 740 × 8740 RDMs) yielded parcel-wise results that were almost perfectly correlated with our session-wise computation (Pearson’s $r=0.995$). While the full-set computation produced systematically lower alignment values (linear slope $\approx 0.53$; $R^{2}=0.99$), this correlation validates that our session-wise approach preserves the relative spatial pattern of the findings.
B.2 BOLD5000 dataset
@@
We processed the BOLD5000 dataset 24 using the same session-wise pipeline as for THINGS-fMRI. This dataset comprised 3T recordings from four participants performing a valence rating task on 4,916 unique images drawn from the MS-COCO, ImageNet, and SUN databases. We used the pre-processed, single-trial $\beta$-estimates provided by the authors, selecting only the first presentation of each image. Inter-subject and model-brain RSA were computed following the same session-wise procedure described above.
@@
For voxel extraction, we registered the HCP-MMP1.0 atlas from MNI152 standard space to each subject’s native T1-weighted scan and then to their mean functional volume, using FSL’s FLIRT and FNIRT. While qualitative inspection confirmed consistent parcel definitions across participants, this volume-to-volume registration has a lower effective resolution compared to the specialized registration available for the NSD atlas. This may account for some parcel-specific results, such as the lower-than-expected alignment values observed in certain regions (e.g., MST). All other preprocessing and computational steps paralleled the main NSD pipeline.
C Implementation details for large-scale RSA comparisons
@@
To efficiently handle the computational demands associated with large-scale RSA analyses (due to dataset size and permutation complexity), we developed a GPU-accelerated implementation using PyTorch84. Specifically, we reformulated representational dissimilarity matrix (RDM) calculations as optimized matrix operations. For each participant pair $p,q\in\mathcal{P}$, and each region $r\in\mathcal{R}$, we constructed stacked matrices consisting of flattened RDMs across all cortical regions, significantly reducing computational overhead and runtime compared to conventional RSA implementations.
@@
Given data matrices for matched trials $X_{p,r}[\boldsymbol{i}_{pq},:]$ and $X_{q,r^{\prime}}[\boldsymbol{j}_{pq},:]$, we computed upper-triangular vectors of dissimilarities:
$\mathbf{x}^{(p)}_{pq,r}=\operatorname{vec}_{u}D\left(X_{p,r}[\boldsymbol{i}_{pq},:]\right),\quad\mathbf{x}^{(q)}_{pq,r^{\prime}}=\operatorname{vec}_{u}D\left(X_{q,r^{\prime}}[\boldsymbol{j}_{pq},:]\right),$ (C5)
@@
where $D(\cdot)$ denotes the dissimilarity matrix and $\operatorname{vec}_{u}$ denotes vectorization of the upper triangle.
@@
Each RDM vector was centered and $l_{2}$-normalized before stacking into a matrix containing all ROIs:
@@
$\mathbf{X}^{(p)}_{pq}=\begin{bmatrix}\tilde{\mathbf{x}}^{(p)}_{pq,1}\\ \tilde{\mathbf{x}}^{(p)}_{pq,2}\\ \vdots\\ \tilde{\mathbf{x}}^{(p)}_{pq,|\mathcal{R}|}\end{bmatrix}\quad\in\mathbb{R}^{|\mathcal{R}|\times\frac{n(n-1)}{2}},$ (C6)
@@
The inter-subject connectivity matrix was then obtained as:
$\mathcal{A}^{\mathrm{IS}}_{p,q}[r,r^{\prime}]=\mathbf{X}^{(p)}_{pq}\left(\mathbf{X}^{(q)}_{pq}\right)^{\top}\in\mathbb{R}^{|\mathcal{R}|\times|\mathcal{R}|}$ (C7)
Analogous formulations were applied for Spearman-based RSA, with rank-normalization applied to RDM vectors prior to computing dot products.
For brain–model alignment, similar stacked matrices were constructed separately for brain RDMs and model RDMs across sessions. The brain matrix was organized as:
$\mathbf{X}^{(\text{brain})}_{p}\in\mathbb{R}^{|\mathcal{R}|\times|\mathcal{T}_{p}|\times\frac{n(n-1)}{2}},$ (C8)
and the model matrix as:
$\mathbf{X}^{(\text{model})}_{m}\in\mathbb{R}^{|\mathcal{L}_{m}|\times|\mathcal{T}_{p}|\times\frac{n(n-1)}{2}}.$ (C9)
Brain–model RSA alignments across regions, layers, and sessions were computed as the inner product between normalized RDM vectors for each session:
$\mathcal{C}_{p,r,m,s}(l)=\langle\mathbf{X}^{(\text{brain})}_{p,r,s},\;\mathbf{X}^{(\text{model})}_{m,l,s}\rangle.$ (C10)
This operation was implemented as a contraction corresponding to the einsum notation $r\,s\,d,l\,s\,d\to r\,l$.
For permutation testing, we permuted stimulus labels defined by the permutation $\sigma\in S_{n}$ prior to RSA computation. Rather than recomputing dissimilarity matrices after each permutation, we constructed an equivalent permutation $\sigma^{\prime}$ over the vectorized upper-triangular RDM entries, matching the induced pairwise index shifts. This allowed permutation of the flattened RDMs directly via:
$\mathbf{X}^{(p)}_{pq}[:,\sigma^{\prime}],$ (C11)
@@
D Extended figures
Figure D1: Detailed parcel-level alignment for all modalities. Box plots show representational alignment scores computed for each of the 180 cortical parcels of the symmetric HCP atlas 31, organized by macro-anatomical groups 47. Each box represents the distribution of alignment scores across the eight participants ($N=8$). The red line and shaded area in each panel denote the mean and standard deviation of the null distribution, respectively, estimated via permutation testing. (A) Inter-subject alignment (IS-RSA). (B) Brain-to-vision-model alignment. (C) Brain-to-language-model alignment. This figure provides a detailed view of the results summarized in the main text, showing the parcel-by-parcel variability and confirming the concentration of high alignment within the Early Visual, Ventral, and LOTC hubs.
Figure D2: Cortical flat map projections of representational alignment. To provide a comprehensive view of their spatial distribution, group-level alignment scores are projected onto a flattened cortical surface. The figure shows (A) inter-subject alignment, (B) vision-model alignment, and (C) language-model alignment. Values represent group-averaged RSA scores (Pearson’s $\rho$, $N=8$ from the NSD dataset), and major sulci are labeled for anatomical orientation. This visualization makes the full extent of the alignment patterns clear, particularly highlighting the widespread negative alignment (cool colors) between language models and the ventral visual stream, in contrast to the positive alignment seen for inter-subject and vision-model comparisons.
Figure D3: Detailed layer-wise alignment profiles across cortical parcels. This figure provides a detailed, parcel-by-parcel view of the brain-model alignment data summarized in the main text (Fig. 3), showing results for the 15 macro-anatomical groups with the highest inter-subject alignment. Only parcels that were themselves statistically significant in the inter-subject analysis are displayed. For each modality, the plotted alignment curves represent the RSA score averaged across all models of that type. Each individual line shows the mean alignment for a single parcel across participants ($N=8$), with shaded areas indicating the SEM. (A) Brain-to-vision-model alignment. This detailed view confirms that the distinct alignment profiles—decreasing, distributed, and increasing—are characteristic of the main hubs and often extend to anatomically adjacent parcels (e.g., the increasing profile of the LOTC hub is also observed in nearby STS regions). In contrast, areas with lower overall correspondence, such as in the prefrontal cortex, tend to show weak alignment across all model layers without a clear hierarchical preference. (B) Brain-to-language-model alignment. This view confirms that positive alignment is restricted to the LOTC hub and a small number of anatomically proximal parcels, all of which consistently exhibit the same step-like alignment profile.
Figure D4: Brain-model alignment profiles for different model families. Layer-wise alignment (RSA, Pearson’s $r$) is shown for representative parcels from the three main hubs, computed separately for each family of models (see Table 1 for a full list). Each curve represents the mean alignment across participants ($N=8$), with shaded areas indicating the SEM. Vision Models: Most vision model families (AugReg, CLIP, DINOv2) reproduce the characteristic hierarchical profiles: a decreasing alignment with Early Visual Cortex, a distributed alignment with the Ventral Hub, and an increasing alignment with the LOTC Hub. The Masked Autoencoder (MAE) models are a notable exception; consistent with their reconstructive training objective, they maintain a high alignment with Early Visual Cortex across all layers. Language Models: Despite differences in architecture, training data, and fine-tuning (e.g., instruction tuning in BLOOMZ), all language model families exhibit a similar pattern. Alignment is negligible or negative for Early Visual and Ventral hub parcels, while all families show the same characteristic step-function profile for parcels in the LOTC hub.
Figure D5: Comparison of within-subject and inter-subject representational analyses. (A) Cortical map of within-subject alignment (RSA, Pearson’s $r$). This was computed analogously to the inter-subject analysis by comparing RDMs from different trial repetitions within each participant, and then averaging across participants ($N=8$). The map reveals a similar spatial distribution to the inter-subject version, with a posterior-to-anterior gradient and three distinct hubs. (B) Parcel-wise comparison of within-subject and inter-subject alignment. The two measures are tightly correlated across the cortex, following a power-law relationship (fit shown in red; $R^{2}=0.98$). This indicates that brain regions that are consistent within an individual are also the ones that are consistent across individuals. (C) Within-subject representational connectivity matrix. Computed by correlating RDMs between all pairs of parcels within each participant, this analysis reveals the same three-hub block structure discovered using the inter-subject measurements. (D) Directed connectivity graph of the within-subject data, pruned using a 2-minimum spanning tree to highlight the network backbone. Edge color and width encode the inter-region alignment (RSA value), while node color represents each parcel’s peak alignment depth with vision models. Directionality is inferred from this depth measure based on a low-to-high sequential hierarchy, with arrows pointing from shallower- to deeper-aligning regions. The graph confirms the same information flow as the inter-subject analysis, with two streams emerging from early visual cortex: a medial-ventral stream to the Ventral hub and a lateral-dorsal stream to the LOTC hub.
Figure D6: Whole-cortex representational connectivity network. Representational connectivity network among the 157 cortical parcels that exhibited statistically significant inter-subject alignment. The network is pruned using a 2-minimum spanning tree to highlight the strongest connections. To minimize confounds from anatomical variability, connectivity was computed on a within-subject basis ($N=8$); for each participant, we correlated the RDM of every parcel with every other parcel across different repetitions of the same stimuli, then averaged the resulting connectivity matrices. Edge color encodes the strength of this inter-region alignment (RSA), while edge width indicates the spanning tree order (1st or 2nd) to highlight the most central pathways. Node color represents each parcel’s peak alignment depth with vision models, providing a proxy for its hierarchical position. Arrows are added to the pruned edges to denote a statistically significant difference in this depth between connected parcels (paired $t$-test, $t(7)$, FDR-corrected $p<0.05$), indicating the putative direction of information flow. The graph confirms the two primary visual processing streams identified in the main text. A medial-ventral stream connects early visual areas to core ventral stream regions involved in scene and object processing. In parallel, a lateral-dorsal stream connects early visual areas to the LOTC hub, including motion-sensitive areas (MT+ complex) and higher-order regions in the superior temporal sulcus (STS) and temporoparietal junction (TPOJ). The analysis also reveals key bridges between these streams, with nodes such as LO3 and PGp linking the ventral and LOTC hubs. Furthermore, despite weaker overall alignment, frontal and parietal regions like the Frontal Eye Fields (FEF) and the Medial Intraparietal Area (MIP) emerge as distinct hubs, suggesting their integration into this large-scale, stimulus-driven network.
Figure D7: Biological content selectively drives LOTC alignment and lateral-stream connectivity. Scenes were split into those that contained biological agents (people or animals; 65.3% of the shared image set) and those that did not, and all inter-subject analyses were recomputed on each subset. (A) Parcel-wise inter-subject RSA within the three hubs. Blue: scenes with biological agents; orange: scenes without them. Early Visual parcels (V1–V4) show no reliable difference, Ventral parcels (VMV1–PHA3) align more for non-biological scenes, whereas LOTC parcels (V4t–TPOJ3) align almost exclusively for biological scenes. Stars mark two-tailed paired $t$-tests across eight participants (${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$, FDR-corrected). (B,C) Group-level representational-connectivity matrices (left) and backbone graphs (right) for scenes without (B) and with (C) biological agents. Backbones were extracted with a three-iteration minimum-spanning-tree procedure. Node colour encodes each parcel’s peak vision-model layer (early $\rightarrow$ late: green $\rightarrow$ purple), node size scales with within-parcel IS-RSA, edge hue shows pairwise RSA, and edge width denotes the iteration (1–3) at which the edge entered the spanning tree. The lateral stream linking Early Visual Cortex to LOTC is largely absent for non-biological scenes (B) but re-emerges when biological agents are present (C), mirroring the parcel-level results in (A).
Figure D8: Image-based atlas of the dominant representational dimensions in Early Visual Cortex. The first two shared representational components for the Early Visual Cortex hub, derived using KMCCA on the responses of all participants to the shared image set. Each point is an individual stimulus image, and its colored border indicates its general semantic category (see legend). The layout reveals no clear clustering by semantic category. Instead, images that are close to each other in this shared space often have similar low-level visual properties (e.g., color palettes, textures, or global shapes), confirming that the geometry of this hub is primarily driven by visual similarity rather than abstract content.
Figure D9: Image-based atlas of the dominant representational dimensions in the Ventral Hub. The first two shared representational components for the Ventral hub, derived using KMCCA. Each point is an individual stimulus image. The layout reveals an organizing principle: a smooth gradient from scene-dominated images to object-dominated images. For example, the upper-left quadrant contains images of indoor spaces, the lower portion contains outdoor scenes, and the upper-right quadrant contains images where a single object is the primary focus. This evidences that the dominant representational axis of this hub recapitulates the classic scene-to-object gradient of ventromedial cortex.
Figure D10: Image-based atlas of the dominant representational dimensions in the LOTC Hub. The first two shared representational components for the LOTC hub, derived using KMCCA. Each point is an individual stimulus image, with its border colored by its general semantic category (see legend). The first KMCCA component (the x-axis) separates images containing biological agents (left cluster) from those without (right cluster). Within the biological cluster, there is a further separation, with images of animals generally located to the right of the images of people. This confirms that the LOTC hub’s geometry is primarily organized around the presence and type of animate entities in a scene.

### v2

@@ Acknowledgments
Figure 17: Global power-law model captures pairwise relations among all RSA metrics. Scatter plots (gray points) show parcel-wise observed RSA values for all metric combinations: inter-subject (shifted and unshifted), within-subject, subject-versus-group, and vision-model correlations, each additionally split by hemisphere sampling (left, right, both). Red curves depict an independent power-law fit for that specific scatter, whereas black dashed curves are the predictions from a single joint fit of Eq. 2 applied to all metrics simultaneously (no panel-specific re-fitting). The accompanying table reports the fitted scale factors $a_{j}$, exponents $b_{j}$, and explained variance $R^{2}$ for every metric (model-predicted vs observed RSA). All metric trends are explained with $R^{2}>0.92$, confirming that one latent similarity map, together with metric-specific attenuation parameters, suffices to account for the entire measurement set.
Supplementary Note 5: Token‐occurrence control for language–brain RSA
To quantify how much of the language–brain alignment reflects initial tokenization, we constructed RDMs from token-occurrence vectors. Each caption was encoded as a count vector—each entry recording the number of times a given token appears—and pairwise Pearson dissimilarities among these vectors yielded a tokenizer-based RDM. We applied this procedure to six vocabularies (BLOOMZ, Gemma-2, LLaMA-1/2/3; see main text Table 1), covering all tokenizers used in our language models, and then computed parcel-wise RSA against NSD fMRI responses using the symmetric HCP atlas. Parcel-wise alignments were nearly identical across tokenizers (pairwise Spearman’s $\rho\geq 0.978$), so we report their average.
The tokenizer–brain alignment map closely matches that of full language models: it peaks in the LOTC hub and shows minimal or negative alignment in early visual parcels (Suppl. Fig. 18A,D). Token-count geometry correlates strongly with each model’s first-layer alignment (Pearson’s $\rho=0.78$; Suppl. Fig. 18B) and with maximal alignment across layers ($\rho=0.68$; Suppl. Fig. 18C), indicating that token frequencies account for a substantial portion of the language–brain correspondence.
Repeating the analysis with a word tokenizer—splitting captions on whitespace rather than into subword units—again yielded pronounced LOTC alignment (Suppl. Fig. 18E). These findings suggest that LOTC’s alignment to language models largely reflects sensitivity to discrete lexical tokens common across descriptions—functioning as a “bag-of-words” representation—rather than to compositional or syntactic structure 100.
Figure 18: Tokenizer‐level contributions to language–brain alignment. We derived RDMs from token‐occurrence count vectors over each model’s vocabulary and computed parcel‐wise RSA against NSD fMRI responses (averaged across the six tokenizers used in our language models). (A) Scatter of tokenizer‐brain vs. inter‐subject alignment, with peak values in the LOTC hub. (B) Token‐count alignment against each model’s first‐layer alignment. (C) Token‐count alignment against each model’s maximal alignment across layers. (D) Cortical map of tokenizer‐brain alignment, showing pronounced LOTC localization. (E) Alignment map for a whitespace‐based word tokenizer, which likewise peaks in LOTC, consistent with a “bag‐of‐words” explanation rather than richer compositional structure.
Supplementary Note 6: Robustness to Similarity Metric (Spearman‐RSA & CKA)
Figure 19: Comparison of alignment metrics. (A–C) Spearman‐RSA. Replacing Pearson’s $r$ with Spearman’s rank correlation yields virtually identical maps: the inter‐subject alignment (A), the parcel‐wise scatter against vision‐ and language‐model alignment (B), and the representational connectivity matrix (C) all match the Pearson‐RSA findings, with parcel‐wise strengths correlating at $r=0.99$. (D–F) Unbiased CKA. Centered Kernel Alignment, which captures global dependence via an unbiased HSIC estimator, again highlights Early Visual, Ventral, and LOTC hubs (D). In the parcel‐wise comparison (E), CKA inflates LOTC alignment and attenuates early visual alignment–reflecting its sensitivity to low‐dimensional, dominant components–yet CKA and Pearson‐RSA remain strongly correlated ($r=0.91$). The CKA‐based connectivity matrix (F) preserves the two‐stream topology, though as a dependence measure it does not distinguish signed similarity.
Recent work has cautioned that representational similarity findings can depend not only on stimulus design but also on the choice of similarity metric—some measures probe local, pairwise geometry while others emphasize global, space-wide dependence 25. To test the robustness of our results to these methodological choices, we repeated key analyses using two alternatives to Pearson-RSA: (i) rank-based Spearman-RSA, which preserves only the ordering of pairwise dissimilarities, and (ii) Centered Kernel Alignment (CKA), which quantifies global statistical dependence via an unbiased Hilbert-Schmidt Independence Criterion (HSIC) estimator 83.
Replacing Pearson with Spearman correlation yields a purely rank-based, nonparametric measure agnostic to absolute distance magnitudes. As shown in Supplementary Fig. 19A–C, the inter-subject map, the parcel-by-parcel scatter against model alignment, and the representational connectivity matrix all mirror our original Pearson-RSA results (parcel-wise Spearman’s $\rho=0.99$). This confirms that neither the linearity nor the distributional assumptions of Pearson correlation dictate our findings—our local pairwise geometry holds under a fully nonparametric test.
CKA measures dependence across entire representational spaces, prioritizing alignment along principal axes rather than isolated pairwise distances. We computed linear-kernel CKA (using the unbiased HSIC estimator 83) under the same shifted-repetition protocol. Supplementary Fig. 19D–F shows that CKA again recovers the three hubs but inflates LOTC alignment relative to Early Visual cortex in the parcel scatter. This reflects CKA’s known sensitivity to low-dimensional, dominant components (often stronger in higher-level cortex) over the high-dimensional, finer-grained structure of early vision. Despite this spectral re-weighting, parcel-wise CKA and Pearson-RSA correlate strongly ($r=0.91$), and the CKA connectivity matrix preserves the core two-stream topology, even though CKA is strictly non-negative and thus cannot index anticorrelated geometries.
Supplementary Note 7: Shared Component Decomposition and Partial RSA Controls
Figure 20: Extended analysis of representational dimensions. (A) KMCCA projections onto the top two canonical axes for six representative parcels: V1, V4, VMV2, PHA2, MT, and TPOJ2. V1 shows no clear clustering; V4 begins to separate by content; Ventral and LOTC parcels replicate the hub-level semantic gradients. (B) Words ranked by loading on the first KMCCA component (positive loadings in blue, negative in red) for each hub. Early Visual words share low-level shape features; Ventral words span scene vs. object terms; LOTC words distinguish animate vs. inanimate labels. (C) Partial RSA boxplots for the three hubs, showing inter-subject alignment under no control (baseline), controlling for the scene-object gradient, or biological presence. LOTC alignment is markedly reduced by each control, whereas Early Visual and Ventral hubs are minimally affected.
This supplementary note details our use of Kernel Multi-view Canonical Correlation Analysis (KMCCA) 40 to isolate the axes driving our RSA findings. While RSA provides a single global similarity score, KMCCA decomposes this correspondence into a ranked set of orthogonal “representational channels,” revealing which latent axes contribute most strongly to the shared geometry.
We performed KMCCA using the Python package mvlearn (v0.5.0) 68 with a correlation kernel and a regularization parameter of $\lambda=0.3$, utilizing the implementation’s default parameters. For each cortical hub, we formed a data matrix for each participant from the GLM-denoised single-trial $\beta$-estimates. We included only those trials for which each (image, repetition) pair was available for all eight participants. This procedure yielded a set of eight matrices ($X_{p,r}\in\mathbb{R}^{m\times v_{p,r}}$), one for each participant, all representing an identical sequence of stimulus trials. The KMCCA algorithm was then applied to this set of matrices, returning projections onto a common shared subspace. For visualization (Supplementary Fig. 20A; Fig. 5A-C), we projected each subject’s data onto the first two canonical axes and then averaged the projections for the repetitions of each unique image to produce a single, stable point per stimulus.
The parcel-level KMCCA projections (Supplementary Fig. 20A) confirmed that the organizational principles observed at the hub level were consistent within their constituent parcels. In Early Visual Cortex, the V1 parcel showed no clear semantic organization in its first two dimensions, whereas the V4 parcel already exhibited emergent clustering by image content. Within the Ventral hub, representative parcels such as VMV2 and PHA2 reproduced the hub-level scene-to-object gradient and showed clear organization by semantic category. Similarly, within the LOTC hub, parcels like MT and TPOJ2 showed the same primary separation between biological and non-biological stimuli. Within these main clusters, stimuli were further organized by fine-grained semantic categories, mirroring the main hub-level analysis (Fig. 5C).
To interpret these KMCCA axes semantically, we projected the image captions onto the shared subspace learned for each hub. We first created a vocabulary from the union of all words present in the captions ($s=11,090$ unique words), encoding each caption as a binary word-occurrence vector (a “bag-of-words” representation). We then projected this caption matrix onto the first KMCCA component derived for each hub, yielding a loading weight for each word in the vocabulary. By ranking these weights, we identified the words that contributed most positively and negatively to each dimension (Fig. 20B). This word-loading analysis confirmed our interpretation of each hub’s organizing principle: in Early Visual Cortex, top-weighted words were not semantically coherent but appeared to share coarse visual features; in the Ventral hub, positive tokens index scene contexts (e.g., ‘kitchen’, ‘bathroom’) while negative tokens name isolated objects (e.g., ‘plate’, ‘frisbee’); and in the LOTC hub, tokens cleanly separated animate from inanimate labels. This analysis provides a semantic interpretation for the organization of each cluster, which can be visually confirmed in the corresponding image atlases (Supplementary Figs. 911).
Finally, we quantified the impact of the identified dimensions on the overall inter-subject alignment using partial RSA. Denoting $\rho(X,Y)$ as the standard Pearson correlation used to compute the alignment between two RDMs, and $Z$ as a control RDM (e.g., one built from the first KMCCA component or from a binary “biological presence” vector), the partial RSA was computed as:
$\mathrm{pRSA}(X,Y\mid Z)=\frac{\rho(X,Y)\;-\;\rho(X,Z)\,\rho(Y,Z)}{\sqrt{(1-\rho(X,Z)^{2})\,(1-\rho(Y,Z)^{2})}}.$ (3)
As shown in the main text (Fig. 5d-f), controlling for the continuous, data-driven KMCCA axis produced only minimal drops in the Early Visual and Ventral hubs, but a substantial reduction in the LOTC hub. A complementary analysis confirmed this result: controlling for the discrete categorical variables yielded a drop in alignment of a similar magnitude to that produced by the continuous KMCCA components (Supplementary Fig. 20C).
Supplementary Note 8: THINGS-fMRI and BOLD5000 Processing
THINGS-fMRI dataset
@@
We processed the THINGS-fMRI dataset 46 using a pipeline analogous to that for NSD. This dataset consists of 3T fMRI recordings from three participants performing an oddball detection task on 8,740 unique object images. A subset of 100 images was repeated once per session (12 presentations total); all other images appeared only once. For our primary computations, we used only the first presentation of each of the 8,740 images.
@@
$\mathcal{A}_{p,q,r,r^{\prime},s}^{\mathrm{IS}}=\mathrm{RSA}\bigl(X_{p,r,s},X_{q,r^{\prime},s}\bigr).$ (4)
@@
Group-level maps were then obtained by averaging these scores across sessions and all participant pairs. Model-brain RSA followed the same session-wise procedure. As a control, we confirmed that computing IS-RSA over the full 8,740-trial set (using $8,740\times 8,740$ RDMs) yielded parcel-wise results that were almost perfectly correlated with our session-wise computation (Pearson’s $r=0.995$). While the full-set computation produced systematically lower alignment values (linear slope $\approx 0.53$; $R^{2}=0.99$), this correlation validates that our session-wise approach preserves the relative spatial pattern of the findings.
BOLD5000 dataset
@@
We processed the BOLD5000 dataset 11 using the same session-wise pipeline as for THINGS-fMRI. This dataset consists of 3T recordings from four participants performing a valence rating task on 4,916 unique images drawn from the MS-COCO, ImageNet, and SUN databases. We used the pre-processed, single-trial $\beta$-estimates provided by the authors, selecting only the first presentation of each image. Inter-subject and model-brain RSA were computed following the same session-wise procedure described above.
@@
For voxel extraction, we registered the HCP-MMP1.0 atlas from MNI152 standard space to each subject’s native T1-weighted scan and then to their mean functional volume using FSL’s FLIRT and FNIRT. Qualitative inspection confirmed consistent parcel definitions across participants. All other preprocessing and computational steps paralleled the main NSD pipeline.
Supplementary Note 9: Implementation details for large-scale RSA comparisons
@@
To efficiently handle the computational demands of large-scale RSA analyses–driven by dataset size and extensive permutation testing–we developed a GPU-accelerated implementation using PyTorch. Specifically, we reformulated representational dissimilarity matrix (RDM) calculations as optimized batch matrix operations. For each participant pair $p,q\in\mathcal{P}$ and cortical region $r\in\mathcal{R}$, we constructed stacked matrices containing flattened RDMs for all regions simultaneously, significantly reducing memory overhead and runtime compared to iterative loop-based implementations.
@@
Given data matrices for matched trials $X_{p,r}[\boldsymbol{i}_{pq},:]$ and $X_{q,r^{\prime}}[\boldsymbol{j}_{pq},:]$ (where the number of matched trials is $N_{\text{trials}}$), we first computed the upper-triangular vectors of dissimilarities:
$\mathbf{x}^{(p)}_{pq,r}=\operatorname{vec}_{u}D\left(X_{p,r}[\boldsymbol{i}_{pq},:]\right),\quad\mathbf{x}^{(q)}_{pq,r^{\prime}}=\operatorname{vec}_{u}D\left(X_{q,r^{\prime}}[\boldsymbol{j}_{pq},:]\right),$ (5)
@@
where $D(\cdot)$ denotes the dissimilarity function (e.g., 1Pearson r1-$\text${Pearson }r) and $\operatorname{vec}_{u}$ denotes vectorization of the upper triangle.
@@
Each RDM vector was centered and $l_{2}$-normalized before being stacked into a single matrix containing all ROIs:
@@
$\mathbf{X}^{(p)}_{pq}=\begin{bmatrix}\tilde{\mathbf{x}}^{(p)}_{pq,1}\\ \tilde{\mathbf{x}}^{(p)}_{pq,2}\\ \vdots\\ \tilde{\mathbf{x}}^{(p)}_{pq,|\mathcal{R}|}\end{bmatrix}\quad\in\mathbb{R}^{|\mathcal{R}|\times\frac{N_{\text{trials}}(N_{\text{trials}}-1)}{2}},$ (6)
@@
The full inter-subject connectivity matrix was then obtained efficiently via a single matrix multiplication:
$\mathcal{A}^{\mathrm{IS}}_{p,q}=\mathbf{X}^{(p)}_{pq}\left(\mathbf{X}^{(q)}_{pq}\right)^{\top}\in\mathbb{R}^{|\mathcal{R}|\times|\mathcal{R}|}$. (7)
Analogous formulations were applied for Spearman-based RSA, with rank-transformation applied to RDM vectors prior to normalization and dot products.
For brain–model alignment, we constructed similar stacked tensors for brain and model RDMs across sessions. The brain tensor was organized as:
$\mathbf{X}^{(\text{brain})}_{p}\in\mathbb{R}^{|\mathcal{R}|\times S\times\frac{N_{\text{trials}}(N_{\text{trials}}-1)}{2}},$ (8)
and the model tensor as:
$\mathbf{X}^{(\text{model})}_{m}\in\mathbb{R}^{L\times S\times\frac{N_{\text{trials}}(N_{\text{trials}}-1)}{2}},$ (9)
where $S$ is the number of sessions and $L$ is the number of model layers.
Brain–model RSA alignments across all regions and layers were computed via a batched contraction of the normalized RDM vectors:
$\mathcal{C}_{p,m}=\langle\mathbf{X}^{(\text{brain})}_{p},\;\mathbf{X}^{(\text{model})}_{m}\rangle$. (10)
This operation was implemented as an Einstein summation corresponding to the signature $r\,s\,d,l\,s\,d\to r\,l$ (summing over sessions $s$ and dissimilarity pairs $d$).
For permutation testing, we permuted stimulus labels according to $\sigma\in S_{N_{\text{trials}}}$ prior to RSA computation. Rather than recomputing dissimilarity matrices from raw data after each shuffle (an $O(N_{\text{trials}}^{2})$ operation), we pre-calculated the equivalent permutation $\sigma^{\prime}$ over the vectorized upper-triangular indices that matches the induced pairwise index shifts. This allowed us to permute the flattened RDMs directly via:
$\mathbf{X}^{(p)}_{pq}[:,\sigma^{\prime}],$ (11)